OpenStudy (anonymous):

Find the imaginary or real solutions of the equation by factoring X^3+64=0

pooja195 (pooja195):

Hi :) Welcome to openstudy! Any ideas on what the first step maybe?

OpenStudy (anonymous):

No. I'm not sure how to factor numbers like this one

pooja195 (pooja195):

Ok well similar to difference of 2 squares we can start by subtract 64 from both sides \[\huge~\rm~x^3+64-63=0 -64\]

pooja195 (pooja195):

Now we have \[\huge~\rm~\bf~x^3=-64\] Now you would take the cube root are you familiar with how to do that?

OpenStudy (anonymous):

Of I had a calculator

pooja195 (pooja195):

Or without a calculator

zepdrix (zepdrix):

That's one good approach c: It looks like they want us to solve this one by factoring though. Recall your `Sum of Cubes` Formula:\[\large\rm a^3+b^3=(a+b)(a^2-ab+b^2)\]

pooja195 (pooja195):

I didn't see they wanted it by factoring sorry im not familiar with that method @zepdrix seems to be good at it perhaps he can continue with the problem :-)

zepdrix (zepdrix):

64 is actually a perfect cube, 64 = 4*4*4 So we can rewrite our expression,\[\large\rm x^3+64\]as\[\large\rm x^3+4^3\]From there, we can apply our `Sum of Cubes` Formula

OpenStudy (anonymous):

Okay, so I find the cube, and then re-write it.

zepdrix (zepdrix):

Applying the formula gives us this,\[\large\rm x^3+4^3=(x+4)(x^2-4x+4^2)\] I'll give you a moment to digest that, look back at the formula involving a and b for reference.

OpenStudy (anonymous):

I don't get how to use the formula. I've never seen it in my life

zepdrix (zepdrix):

Have you ever heard the acronym SOAP? It's often used as a nice easy way to remember this formula.

OpenStudy (anonymous):

No I have not. My math teacher doesn't put any description in his online lessons. He kind of just expects us to understand everything

zepdrix (zepdrix):

This is where our SOAP acronym comes from.